OpenStudy (anonymous):

What is the equation in standard form of the line which passes through (1, −3) and has a slope of 2?

OpenStudy (anonymous):

2x − y = 5 or 2x − y = −5 or 2x + y = 5 or 2x + y = −5

OpenStudy (anonymous):

i chose a

ganeshie8 (ganeshie8):

why

OpenStudy (anonymous):

i did y+3=2x + 2 but after that im confused on what to do

ganeshie8 (ganeshie8):

we're given a point : \((1, -3)\)

ganeshie8 (ganeshie8):

we're also given slope = m = \(2\)

ganeshie8 (ganeshie8):

so, point slope form of line is, \(y - y_1 = m(x-x_1)\)

ganeshie8 (ganeshie8):

\(y--3 = 2(x-1)\)

ganeshie8 (ganeshie8):

\(y+3 = 2x-2\)

OpenStudy (anonymous):

okay yes i got that far

ganeshie8 (ganeshie8):

\(3 = 2x-y-2\)

ganeshie8 (ganeshie8):

\(5 = 2x-y\)

ganeshie8 (ganeshie8):

\(2x-y = 5\)

ganeshie8 (ganeshie8):

So you're right !

OpenStudy (anonymous):

yay